OpenStudy (anonymous):

Given the point (5, 6) and the slope of 5, find y when x = 21. 86 90 105 111

OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (5, 7) and (8, 22)? y = −5x + 18 y = 5x − 18 y = −5x − 18 y = 5x + 18

OpenStudy (anonymous):

(5,6) means when x=5 then y=6 The slope tells you how much does y increase when you increase x by 1 so in this case x goes up by one then y goes up by 5

OpenStudy (anonymous):

difference between 21 and 5 is 16, meaning you need to get y up by 16. If you devide 16 by 5 then you get 3.2 and that is you x value

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

I need the equation

OpenStudy (anonymous):

aah sec

OpenStudy (anonymous):

ow x=21 i misread it

OpenStudy (anonymous):

y = −5x + 18 y = 5x − 18 y = −5x − 18 y = 5x + 18

OpenStudy (anonymous):

well the slope is 5 so you already know its 5x

OpenStudy (anonymous):

so is it adding 18 or sujbtracting it

OpenStudy (anonymous):

to get the constant you look which value y has when x equals 0. so if you go down by 5 x's you go down by 5 y's because 5*5=25 so 6-25= -19

OpenStudy (anonymous):

It should be y=5x-19 or I made a miscalculation